OpenStudy (jazzyfa30):

Solve the system by using substitution: x - y = 5 2x+y = 13 Select one: a. (6,-1) b. (1,6) c. (-6,-1) d. (6,1)

OpenStudy (jazzyfa30):

@PatrickJordon help

OpenStudy (anonymous):

Ok, What you Want to do is plug in each number for "x" and "y", i did that for you and here's what i got:: X-Y=5 (6)-(1)=5 2x+Y=13 2(6)+(1)=13 So that makes your answer D (6,1)

OpenStudy (jazzyfa30):

thank you I don't know why but everyone else was giving me confusing equations

OpenStudy (anonymous):

Substitute The Numbers In For The Letters Much Easier @Luigi0210

OpenStudy (jazzyfa30):

here is another one

OpenStudy (jazzyfa30):

Use the elimination method to solve the system of equations: 2a+3b=12 5a -3b=9 Select one: a. (2,-3) b. (-3,2) c. (3,2) d. (2,3)

OpenStudy (anonymous):

Bah, idk this one haha

OpenStudy (jazzyfa30):

@amistre64 help

OpenStudy (luigi0210):

add them up and it'll make it so much simpler

OpenStudy (amistre64):

substituting in for x and y IS a form of substitution :)

OpenStudy (anonymous):

Ok its C i missed typed it: 2a+3b=12 2(3)+3(2)=12 AND 5a-3b=9 5(3)-3(2)=9

OpenStudy (jazzyfa30):

but this is elimination method................I'm confused

OpenStudy (anonymous):

dunno if thats right but i tried xD

OpenStudy (jazzyfa30):

ok

OpenStudy (luigi0210):

2a+3b=12 5a -3b=9 ------------------------------ 7a=21 a=3 ------------------------------ 2(3)+3b=12 6+3b=12 3b=6 b=2 -----------------------------------

OpenStudy (anonymous):

# I was Right :)